# Feedback-driven recurrent quantum neural network universality

### v1

@@ 2.2 RQNN approximation of state-space maps and their derivatives
where $C_{j}=\|\widehat{F_{j}}\|_{1}^{2}+4\pi^{2}\|\widehat{F_{j}}\|_{1}\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\sum_{i=1}^{N+d}\xi_{i}^{2}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}$.
@@
Let $j\in\{1,\ldots,N\}$ be fixed. As in the proof of Proposition 2 in [GJ25], we may use the Fourier inversion theorem to represent
@@
Corollary 2.3.
Consider the setting of Proposition 2.2 and assume, in addition, that
$I_{q}=\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\|\bm{\xi}\|^{q}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}<\infty$
for some $q\geq 2$. Then, for any $n\in\mathbb{N}$, there exists $\bm{\theta}\in\bm{\Theta}$ such that for any $j\in\{1,\ldots,N\}$,
where $\bar{C}_{j}=3C_{j}$. Moreover, we can choose $\bm{\theta}=(\bm{\theta}^{1},\ldots,\bm{\theta}^{N})\in\bm{\Theta}^{N}$ with $\bm{\theta}^{j}=(\bm{a}^{i,j},b^{i,j},\gamma^{i,j})_{i=1,\ldots,n}$ in such a way that
$\|\bm{a}^{i,j}\|\leq 2\pi\left(3n\|\widehat{F_{j}}\|_{1}^{-1}\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\|\bm{\xi}\|^{q}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}\right)^{\frac{1}{q}}$ (18)
for all $i=1,\ldots,n$, $j=1\ldots,N$.
The proof of this corollary requires the following lemma, which extends [Gon24, Lemma 4.10].
Lemma 2.4.
@@
The proof mimics that of in [Gon24, Lemma 4.10] by replacing the use of Markov’s inequality for $q=1$ by the more general version:
@@
Proof of the corollary. The corollary follows by replacing the argument leading to (17) in the proof of Proposition 2.2 by Lemma 2.4 and by noticing that
@@
It follows by combining the proof of Proposition 2.2 with the proof of Theorem 3 in [GJ25]. More specifically, the same proof can be used as for Proposition 2.2, except that we need to replace the $L^{2}(\mu)$ error bounds in (15) and (16) by uniform bounds. For (15), we can follow precisely the proof of Theorem 3 in [GJ25] to obtain
@@
with $C_{j}^{\infty,0}=2(\pi+1)\|\widehat{F_{j}}\|_{1}+8\pi M(N+d)^{\frac{1}{2}}\|\widehat{F_{j}}\|_{1}^{\frac{1}{2}}\left(\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\sum_{i=1}^{N+d}\xi_{i}^{2}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}\right)^{1/2}$. Next, we turn to the derivatives, that is, we aim to estimate $\left\|\partial_{k}\bar{F}^{n,\bm{\theta}}_{R,j}-\partial_{k}F_{j}\right\|_{\infty,M}$. Also in this case, we may proceed as in the proof of Theorem 3 in [GJ25] and apply the same estimates to the random variables $U_{i,(\bm{x},\bm{z})}=W_{i}A_{i,k}\sin(B_{i}+\mathbf{A}_{i}\cdot(\bm{x},\bm{z}))$. Let $\varepsilon_{1},\ldots,\varepsilon_{n}$ be i.i.d. Rademacher random variables independent of $\mathbf{A}=(\mathbf{A}_{1},\ldots,\mathbf{A}_{n})$ and $\mathbf{B}=(B_{1},\ldots,B_{n})$. Symmetrisation and independence then yield
@@
$\mathcal{T}$ :=$:=\{(w_{i}a_{i,k}(b_{i}+\mathbf{a}_{i}\cdot(\bm{x},\bm{z})))_{i=1,\ldots,n}\,\colon\,(\bm{x},\bm{z})\in[-M,M]^{N+d}\},$
$\varrho_{i}(x)$ $:=w_{i}a_{i,k}\sin(\frac{x}{w_{i}a_{i,k}}),\quad x\in\mathbb{R},$
for $i=1,\ldots,n$. Then, using the definitions in the first step, the comparison theorem [LT13, Theorem 4.12] in the second step (note $\varrho_{i}(0)=0$ and $\varrho_{i}$ is $1$-Lipschitz), and standard Rademacher estimates (see, e.g., [Gon23]), we obtain
@@
First, extending the proof of Corollary 4 in [GJ25], we show that $F_{j}$ can be approximated on $\mathcal{X}$ up to error $\frac{\varepsilon}{2}$ in $C^{1}$-norm by a function in $C_{c}^{\infty}(\mathbb{R}^{N}\times\mathbb{R}^{d})$. Indeed, first let $M>0$ be such that $\mathcal{X}\subset[-M,M]^{N+d}$. Then, classical approximation results (see, e.g., [Whi34, Lemma 5]) imply that there exists a smooth function $h\colon\mathbb{R}^{N}\times\mathbb{R}^{d}\to\mathbb{R}$ such that
@@
Without loss of generality we may assume that $h\in C^{\infty}_{c}(\mathbb{R}^{N}\times\mathbb{R}^{d})$. Otherwise, we multiply $h$ with a cutoff function $\psi\in C_{c}^{\infty}(\mathbb{R}^{N}\times\mathbb{R}^{d})$ which is equal to $1$ in an open set $U$ with $\mathcal{X}\subset U$ (see, e.g., [Hör90, Theorem 1.4.1]); thereby preserving (22).

### v2

@@ 4.1 RQNN approximation of state-space maps and their derivatives
for any $j\in\{1,\ldots,N\}$, where $C_{j}=\|\widehat{F_{j}}\|_{1}^{2}+4\pi^{2}\|\widehat{F_{j}}\|_{1}\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\sum_{i=1}^{N+d}\xi_{i}^{2}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}$.
Next, we show that it is also possible to also obtain approximation results for QNNs with bounded network coefficients. The proof is provided in Appendix B.3.
Corollary 4.3.
In the setting of Proposition 4.2, assume, in addition, that $\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\|\bm{\xi}\|^{q}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}<\infty$ for some $q\geq 2$. Then, for any $n\in\mathbb{N}$, there exists $\bm{\theta}\in\bm{\Theta}$ such that for any $j\in\{1,\ldots,N\}$,
where $\bar{C}_{j}=3C_{j}$. Moreover, we can choose $\bm{\theta}=(\bm{\theta}^{1},\ldots,\bm{\theta}^{N})\in\bm{\Theta}^{N}$ with $\bm{\theta}^{j}=(\bm{a}^{i,j},b^{i,j},\gamma^{i,j})_{i=1,\ldots,n}$ in such a way that for all $i=1,\ldots,n$, $j=1\ldots,N$,
$\|\bm{a}^{i,j}\|\leq 2\pi\left(3n\|\widehat{F_{j}}\|_{1}^{-1}\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\|\bm{\xi}\|^{q}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}\right)^{\frac{1}{q}}.$ (8)
Proposition 4.4.
The proof can be found in Appendix B.4. Finally, we obtain a qualitative universal approximation result for QNNs jointly with their derivatives. The proof can be found in Appendix B.5.
Corollary 4.5.
4.2 Recurrent QNN approximation bounds for state-space filters
The results in the previous section show that the family of RQNNs that were introduced in (3) is capable of approximating arbitrarily well the very general class of continuously differentiable state-space maps with bounded Fourier transform, together with their derivatives. These approximations hold with respect to both the $L^{2}$ norm (Proposition 4.2 and Corollary 4.3) and the $L^{\infty}$ norm on compacta (Proposition 4.4 and Corollary 4.5). We will now use the uniform RQNN approximation results for the state maps to conclude similar uniform approximation results for the corresponding filters under additional hypotheses that guarantee that those exist.
Consider a state-space system
$\bm{x}_{t}=F(\bm{x}_{t-1},\bm{z}_{t}),\quad t\in\mathbb{Z}_{-},$ (11)
with state process $(\bm{x}_{t})_{t\in\mathbb{Z}_{-}}$ valued in $\mathbb{R}^{N}$, input process $(\bm{z}_{t})_{t\in\mathbb{Z}_{-}}$ valued in $\mathbb{R}^{d}$ and $F\colon\mathbb{R}^{N}\times\mathbb{R}^{d}\to\mathbb{R}^{N}$. We work under the assumption that $F$ is contractive and satisfies Barron-type integrability conditions (Barron, 1992; 1993; Barron & Klusowski, 2018). Then, e.g., Proposition 1 and Remark 2 in Gonon et al. (2020) imply that, for any compact $D_{d}\subset\mathbb{R}^{d}$, the associated filter $U^{F}\colon(D_{d})^{\mathbb{Z}_{-}}\to(B_{N})^{\mathbb{Z}_{-}}$ induced by the restriction of $F$ to $B_{N}\times D_{d}$ is well-defined and continuous.
Our next result shows that among the RQNNs that we discussed in Proposition 4.4 there exist systems that have the echo state property and hence have a filter associated. More importantly, those filters can be used to uniformly approximate any of the filters corresponding to the general systems introduced above in (11) as long as they satisfy a Barron-type integrability condition and are sufficiently contractive. The proof can be found in Appendix C.1. Here, $\|\cdot\|_{2}$ is the spectral norm. In particular, this result shows that the error rate is free from the curse of dimensionality: the error decays as $\frac{1}{\sqrt{n}}$ as we increase $n$, with this rate of decay being independent of the input dimension $d$ and the state space dimension $N$. Thus, the RQNN requires only $\mathcal{O}(\varepsilon^{-2})$ weights and $\mathcal{O}(\lceil\log_{2}(\varepsilon^{-1})\rceil)$ qubits to achieve approximation error $\varepsilon>0$ for the considered state-space systems.
Theorem 4.6.
Suppose $F$ in (11) is continuously differentiable with $\|\nabla_{\bm{x}}F(\bm{x},\bm{z})\|_{2}\leq\lambda$ for all $\bm{x}\in\mathbb{R}^{N},\bm{z}\in D_{d}$ for some $\lambda\in(0,1)$ and, moreover, $F$ satisfies $F_{j}\in\mathcal{F}_{R}$, $\partial_{i}F_{j}\in\mathcal{F}$ and $\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\|\bm{\xi}\|^{4}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}<\infty$ for $j=1,\ldots,N$. Denote by $U^{F}\colon(D_{d})^{\mathbb{Z}_{-}}\to(B_{N})^{\mathbb{Z}_{-}}$ the filter associated to (11). Then for any $n\in\mathbb{N}$ with $n>n_{0}$ there exists $\bm{\theta}\in\bm{\Theta}$ such that the system (4) has the echo state property and the associated filter $\bar{U}\colon(D_{d})^{\mathbb{Z}_{-}}\to(\mathbb{R}^{N})^{\mathbb{Z}_{-}}$ satisfies
Notice that $N$ represents the state space dimension of the target $F$, which is matched by the QRNN dimension to obtain the approximation error bound. Theorem 4.6 also proves an advantage of QRNNs over classical RNNs. RNN approximation bounds for state-space systems driven by Barron-type functions were obtained in (Gonon et al., 2023, Theorem 3). While the approximation rate in Theorem 4.6 is the same ($\frac{1}{2}$ in both cases), the Fourier integrability condition required in the quantum case is strictly weaker. Specifically, the condition $\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\|\bm{\xi}\|^{4}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}<\infty$ implies that the smoothness assumption (Gonon et al., 2023, Definition 2) required for (Gonon et al., 2023, Theorem 3) is satisfied. For example, consider a Sobolev function $F\in H^{s}(\mathbb{R}^{N}\times\mathbb{R}^{d})$. Then, the integrability condition for the QRNN approximation result is satisfied for any $s>\frac{N+d}{2}+4$ (by (Folland, 2020, Lemma 6.5) and its proof). In contrast, the integrability condition for the RNN approximation result in (Gonon et al., 2023, Theorem 3) would require the stronger condition $s>N+d+3$.
4.3 Universality
with $\bm{\theta}=(\bm{\theta}^{1},\ldots,\bm{\theta}^{N})\in\bm{\Theta}^{N}$ and $P_{1},\ldots,P_{N}\in\mathbb{R}^{N\times N}$ linear preprocessing maps. Our modified RQNN is then defined by the state-space system associated to the state map $\tilde{F}^{n,\bm{\theta}}_{R}$
The next lemma (with proof provided in Appendix C.2) shows that adding linear preprocessing maps to reservoir equations can lead to the echo state property without contraction assumptions.
Lemma 4.7.
$\tilde{F}_{j}(\bm{x},\bm{z})=g_{j}(P_{j}\bm{x},\bm{z})$ (15)
Notice that Lemma 4.7 provides the echo state property by imposing a finite memory of K1K-1 time steps on the reservoir. Let $D_{d}\subset\mathbb{R}^{d}$, $B_{m}\subset\mathbb{R}^{m}$ be compact. For a readout $W\in\mathbb{R}^{m\times N}$, denote
$\mathbf{y}_{t}=W\bm{x}_{t}$ (16)
the output process associated to the recurrent QNNs (4) and (14). Our next result proves universality of RQNNs. The proof is provided in Appendix C.3.
Theorem 4.8.
5 Conclusions
Approximation bounds and universality properties are part of the theoretical cornerstone of machine learning models. While some studies have addressed the question of universality for QRC models, the combination of the two had not previously been explored in the context of recurrent QNNs. In this paper, we derived approximation bounds and universality statements for recurrent QNNs based on the circuit implementation presented in Gonon & Jacquier (2025), which is compatible with hardware deployment and whose implementation with Rydberg atoms has been already discussed in Agarwal et al. (2024). This circuit uses a uniformly controlled quantum gate to apply multi-controlled rotations to a set of control and target qubits, and it has been recently shown that it can be efficiently implemented (Zindorf & Bose, 2024; Silva et al., 2024; Zindorf & Bose, 2025).
To prove our results, we first derived approximation bounds for the static version of the QNN and its derivatives. These results are used in Theorem 4.6 to provide filter approximation bounds that show that RQNNs are able to uniformly approximate the filters induced by any contracting Barron-type state-space system. Finally, Theorem 4.8 extends this universality property to the much larger category of arbitrary fading memory, causal, and time-invariant filters. In this last result, neither Barron-type integrability nor contractivity conditions are needed for the target filter. While our results apply to variational systems in which all parameters are trainable, they pave the way for results on quantum reservoir systems in which some parameters in the recurrent layer are randomly generated and only the output layer weights are tuned. Which strategy is best in terms of speed and accuracy will depend on the number of blocks $n$ of the circuit, the intrinsic noise of the hardware, and the target task. Future research will focus on implementing and comparing the variational and reservoir approaches.
This work paves the way for extending the theoretical analysis of QRC models beyond the state-affine system (SAS) paradigm (Martínez-Peña & Ortega, 2023). It is important to understand in which situations the feedback approach is preferable to other protocols. Questions such as the exponential concentration of observables (Sannia et al., 2025; Xiong et al., 2025) and the suitability of QRC models for learning quantum temporal tasks (Tran & Nakajima, 2021; Nokkala, 2023) are fundamental to discerning the conditions that render QRC models more useful than classical machine learning approaches.
While our paper obtains approximation bounds for Barron-type sate-space systems, an important direction of future research will consist in studying approximation error rates for systems with high degrees of roughness or non-contractive dynamics. Furthermore, our paper focuses on approximation properties of RQNNs. Gradient-based training approaches for optimizing RQNN parameters have been proposed, e.g., in Bausch (2020); Li et al. (2023); Siemaszko et al. (2023). Quantum circuit training may face Barren plateaus McClean et al. (2018); Larocca et al. (2025), flat parameter optimization landscapes for large number of qubits. Developing efficient training algorithms and studying these effects in detail will be a further important direction for future research.
Acknowledgments
The authors acknowledge partial financial support from the School of Physical and Mathematical Sciences of the Nanyang Technological University through the SPMS Collaborative Research Award 2023 entitled “Quantum Reservoir Systems for Machine Learning”. RMP acknowledges the QCDI project funded by the Spanish Government. JPO wishes to thank the hospitality of the Donostia International Physics Center and LG and RMP that of the Division of Mathematical Sciences of the Nanyang Technological University, during the academic visits in which some of this work was developed.
Appendix
Appendix A Quantum reservoir computing protocols
For learning problems with temporal structure, quantum reservoir computing (QRC) has emerged as a promising approach for exploiting noisy intermediate-scale quantum (NISQ) technologies. These include ion traps, nuclear magnetic resonance, cold atoms, photonic platforms, and superconducting qubits (Mujal et al., 2021). When implementing QRC models experimentally, it is necessary to consider the backaction and statistical effects introduced by quantum measurements. Backaction refers to the modification of a quantum state after monitoring, also known as wavefunction collapse. Due to the probabilistic nature of quantum theory, measurements must be repeated to compute the expected values of observables, which introduces a statistical component in all these methodologies. Most available experimental implementations rely on the quantum computer paradigm (Dasgupta et al., 2022; Mlika et al., 2023; Suzuki et al., 2022; Yasuda et al., 2023; Chen et al., 2020; Kubota et al., 2023; Molteni et al., 2023; Pfeffer et al., 2022; Ahmed et al., 2025; Hu et al., 2024; Miranda & Shaji, 2025). However, there is an increasing interest in extending this technique to new settings, such as optical pulses (García-Beni et al., 2023; Paparelle et al., 2025), Rydberg atoms (Bravo et al., 2022; Kornjača et al., 2024), and quantum memristors (Spagnolo et al., 2022; Selimović et al., 2025).
Early QRC model implementations relied on the simplest possible approach, namely, the restarting protocol (Dasgupta et al., 2022; Suzuki et al., 2022; Kubota et al., 2023; Chen et al., 2020; Molteni et al., 2023). In this approach, the expected values of observables are obtained by rerunning the algorithm from the first time step at each subsequent time step. This avoids the backaction effect of quantum measurements. However, the complexity of this protocol scales quadratically with the length of the input sequence, making it very time-consuming. A faster alternative is the rewinding protocol (Mujal et al., 2021; Čindrak et al., 2024), where the fading memory of the quantum reservoir is exploited to restart the algorithm with a fixed window of past time steps. This reduces the complexity of the algorithm to linear in terms of input length. Originally proposed in Chen et al. (2020), this protocol has thus far only been considered numerically (Mujal et al., 2023; Čindrak et al., 2024). Both the restarting and rewinding protocols use repetition of previous time steps to reproduce the dynamics of the theoretical model and avoid the disruptive effect of projective measurements used to extract output information. This comes at the cost of halting the quantum dynamics at each time step and the need to buffer the input sequence. Consequently, these approaches lack one of the most important features of traditional reservoir computing, namely, the ability to process information in real time.
New protocols have been proposed to circumvent this problem. The online protocol (Mujal et al., 2023; Franceschetto et al., 2024) uses weak measurements to find a balance between erasing and extracting information. Mid-circuit measurements and reset operations (Hu et al., 2024) can split the reservoir into two parts: memory and readout. The memory retains previous inputs, while measurements only affect the readout part. The feedback protocol (Kobayashi et al., 2024), which can be traced back to QRC with quantum memristors (Spagnolo et al., 2022) and hybrid QRC techniques (Pfeffer et al., 2022; 2023), reinjects the measured observables at each time step as parameters of an input quantum channel. This ensures that no backaction effects are present and that past input information is preserved. Note that in order to compute the observables in real time, these protocols all require several copies of the system to be run in parallel. Furthermore, these protocols can be combined with each other. For instance, the feedback protocol has been combined with both the online protocol (Monomi et al., 2025) and with mid-circuit measurements and reset operations (Murauer et al., 2025).
Of all these approaches, the feedback protocol presents some particularly interesting features. First, the feedback protocol enables us to compute the expected values of observables from a single copy of the system by repeating one time step only. If only a few copies of the system are available, this reduces the experimental time overhead for real-time applications compared to other approaches. Second, in contrast to previous QRC models, where an erasure mechanism is added to provide fundamental properties such as the echo state property, simple unitary operations can provide these properties (Kobayashi et al., 2024). Finally, the dynamical equations of quantum reservoirs under the feedback protocol go beyond the standard state-affine system (SAS) paradigm of QRC models (Martínez-Peña & Ortega, 2023). These properties make the feedback protocol a promising candidate for exploring QRC applications.
Appendix B Proofs for Section 4.1
B.1 Proof of Proposition 4.1
@@ Appendix A Quantum reservoir computing protocols
The proof is a modification of the argument used to obtain (Gonon & Jacquier, 2025, Proposition 1). Recall that
$\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x},\bm{z}):=R-2R[\mathbb{P}_{1}^{n,\bm{\theta}^{j}}(\bm{x},\bm{z})+\mathbb{P}_{2}^{n,\bm{\theta}^{j}}(\bm{x},\bm{z})],\quad(\bm{x},\bm{z})\in\mathbb{R}^{N}\times\mathbb{R}^{d}.$ (18)
Fix $(\bm{x},\bm{z})\in\mathbb{R}^{N}\times\mathbb{R}^{d}$ and $j\in\{1,\ldots,N\}$ and write $\mathbb{P}_{m}:=\mathbb{P}_{m}^{n,\bm{\theta}^{j}}(\bm{x},\bm{z})$ for $m\in\{0,1,2,3\}$. To prove the representation (7), let us first calculate $\mathbb{P}_{m}$.
As a first step, write
[…]
Thus, for $m\in\{0,1,2,3\}$, we have
[…]
$=\frac{1}{n}\sum_{i=0}^{n-1}\left|\left[\mathtt{U}^{(i+1)}_{1}\otimes\mathtt{U}^{(i+1)}_{2}\right]_{m+1,1}\right|^{2}.$
Next, we may calculate
$[\mathtt{U}^{(i)}_{1}\otimes\mathtt{U}^{(i)}_{2}]_{1,1}$ $=[\mathtt{U}^{(i)}_{1}]_{1,1}[\mathtt{U}^{(i)}_{2}]_{1,1}=\cos\left(\frac{\gamma^{i,j}}{2}\right)\cos\left(\frac{b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})}{2}\right),$
$[\mathtt{U}^{(i)}_{1}\otimes\mathtt{U}^{(i)}_{2}]_{2,1}$ $=[\mathtt{U}^{(i)}_{1}]_{1,1}[\mathtt{U}^{(i)}_{2}]_{2,1}=\sin\left(\frac{\gamma^{i,j}}{2}\right)\cos\left(\frac{b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})}{2}\right),$
$[\mathtt{U}^{(i)}_{1}\otimes\mathtt{U}^{(i)}_{2}]_{3,1}$ $=[\mathtt{U}^{(i)}_{1}]_{2,1}[\mathtt{U}^{(i)}_{2}]_{1,1}=\mathrm{i}\cos\left(\frac{\gamma^{i,j}}{2}\right)\sin\left(\frac{b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})}{2}\right),$
$[\mathtt{U}^{(i)}_{1}\otimes\mathtt{U}^{(i)}_{2}]_{4,1}$ $=[\mathtt{U}^{(i)}_{1}]_{2,1}[\mathtt{U}^{(i)}_{2}]_{2,1}=\mathrm{i}\sin\left(\frac{\gamma^{i,j}}{2}\right)\sin\left(\frac{b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})}{2}\right),$
and thus
$\mathbb{P}_{0}$ $=\frac{1}{n}\sum_{i=1}^{n}\cos\left(\frac{\gamma^{i,j}}{2}\right)^{2}\cos\left(\frac{b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})}{2}\right)^{2}$
$\mathbb{P}_{1}$ $=\frac{1}{n}\sum_{i=1}^{n}\sin\left(\frac{\gamma^{i,j}}{2}\right)^{2}\cos\left(\frac{b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})}{2}\right)^{2}$
$\mathbb{P}_{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\cos\left(\frac{\gamma^{i,j}}{2}\right)^{2}\sin\left(\frac{b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})}{2}\right)^{2}$
$\mathbb{P}_{3}$ $=\frac{1}{n}\sum_{i=1}^{n}\sin\left(\frac{\gamma^{i,j}}{2}\right)^{2}\sin\left(\frac{b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})}{2}\right)^{2}.$
Therefore, using $\cos(y)^{2}=\frac{\cos(2y)+1}{2}$, we obtain
$\mathbb{P}_{0}+\mathbb{P}_{1}$ $=\frac{1}{n}\sum_{i=1}^{n}\cos\left(\frac{b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})}{2}\right)^{2}=\frac{1}{2}+\frac{1}{2n}\sum_{i=1}^{n}\cos\left(b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})\right),$
$\mathbb{P}_{0}+\mathbb{P}_{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\cos\left(\frac{\gamma^{i,j}}{2}\right)^{2}=\frac{1}{2}+\frac{1}{2n}\sum_{i=1}^{n}\cos\left(\gamma^{i,j}\right).$
Putting it all together we obtain, for any given $R>0$, that
$\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x},\bm{z})$ $=R-2R[\mathbb{P}_{1}^{n,\bm{\theta}^{j}}(\bm{x},\bm{z})+\mathbb{P}_{2}^{n,\bm{\theta}^{j}}(\bm{x},\bm{z})]$
$=R\left[1+4\mathbb{P}_{0}-2\left(\mathbb{P}_{0}+\mathbb{P}_{1}\right)-2\left(\mathbb{P}_{0}+\mathbb{P}_{2}\right)\right]$
$=\frac{1}{n}\sum_{i=1}^{n}R\cos\left(\gamma^{i,j}\right)\cos\left(b^{i,j}+\bm{a}^{i,j}\cdot(\bm{x},\bm{z})\right).$
B.2 Proof of Proposition 4.2
Proof.
Let $j\in\{1,\ldots,N\}$ be fixed. As in the proof of Proposition 2 in Gonon & Jacquier (2025), we may use the Fourier inversion theorem to represent
@@
B.3 Proof of Corollary 4.3
The proof of this corollary requires the following lemma, which extends Gonon (2024, Lemma 4.10).
Lemma B.1.
@@
The proof mimics that of in Gonon (2024, Lemma 4.10) by replacing the use of Markov’s inequality for $q=1$ by the more general version:
@@
Proof of the corollary. The corollary follows by replacing the argument leading to (28) in the proof of Proposition 4.2 by Lemma B.1 and by noticing that
@@
It follows by combining the proof of Proposition 4.2 with the proof of Theorem 3 in Gonon & Jacquier (2025). More specifically, the same proof can be used as for Proposition 4.2, except that we need to replace the $L^{2}(\mu)$ error bounds in (26) and (27) by uniform bounds. For (26), we can follow precisely the proof of Theorem 3 in Gonon & Jacquier (2025) to obtain
@@
with $C_{j}^{\infty,0}=2(\pi+1)\|\widehat{F_{j}}\|_{1}+8\pi M(N+d)^{\frac{1}{2}}\|\widehat{F_{j}}\|_{1}^{\frac{1}{2}}\left(\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\sum_{i=1}^{N+d}\xi_{i}^{2}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}\right)^{1/2}$. Next, we turn to the derivatives, that is, we aim to estimate $\left\|\partial_{k}\bar{F}^{n,\bm{\theta}}_{R,j}-\partial_{k}F_{j}\right\|_{\infty,M}$. Also in this case, we may proceed as in the proof of Theorem 3 in Gonon & Jacquier (2025) and apply the same estimates to the random variables $U_{i,(\bm{x},\bm{z})}=W_{i}A_{i,k}\sin(B_{i}+\mathbf{A}_{i}\cdot(\bm{x},\bm{z}))$. Let $\varepsilon_{1},\ldots,\varepsilon_{n}$ be i.i.d. Rademacher random variables independent of $\mathbf{A}=(\mathbf{A}_{1},\ldots,\mathbf{A}_{n})$ and $\mathbf{B}=(B_{1},\ldots,B_{n})$. Symmetrisation and independence then yield
@@
$\mathcal{T}$ $=\{(w_{i}a_{i,k}(b_{i}+\mathbf{a}_{i}\cdot(\bm{x},\bm{z})))_{i=1,\ldots,n}\,\colon\,(\bm{x},\bm{z})\in[-M,M]^{N+d}\},$
$\varrho_{i}(x)$ $=w_{i}a_{i,k}\sin(\frac{x}{w_{i}a_{i,k}}),\quad x\in\mathbb{R},$
for $i=1,\ldots,n$. Then, using the definitions in the first step, the comparison theorem (Ledoux & Talagrand, 2013, Theorem 4.12) in the second step (note $\varrho_{i}(0)=0$ and $\varrho_{i}$ is $1$-Lipschitz), and standard Rademacher estimates (see, e.g., Gonon (2023)), we obtain
@@
First, extending the proof of Corollary 4 in Gonon & Jacquier (2025), we show that $F_{j}$ can be approximated on $\mathcal{X}$ up to error $\frac{\varepsilon}{2}$ in $C^{1}$-norm by a function in $C_{c}^{\infty}(\mathbb{R}^{N}\times\mathbb{R}^{d})$. Indeed, first let $M>0$ be such that $\mathcal{X}\subset[-M,M]^{N+d}$. Then, classical approximation results (see, e.g., Whitney, 1934, Lemma 5) imply that there exists a smooth function $h\colon\mathbb{R}^{N}\times\mathbb{R}^{d}\to\mathbb{R}$ such that
@@
Without loss of generality we may assume that $h\in C^{\infty}_{c}(\mathbb{R}^{N}\times\mathbb{R}^{d})$. Otherwise, we multiply $h$ with a cutoff function $\psi\in C_{c}^{\infty}(\mathbb{R}^{N}\times\mathbb{R}^{d})$ which is equal to $1$ in an open set $U$ with $\mathcal{X}\subset U$ (see, e.g., Hörmander, 1990, Theorem 1.4.1); thereby preserving (30).